Suppose you have a $1 million, 5% fixed rate mortgage with annual payments, a maturity of 30 year...

  1. Suppose you have a $1 million, 5% fixed rate mortgage with annual payments, a maturity of 30 years and a balloon payment of $750,000. What is your approximate annual payment?


a. $50,763
b. $51,763

c. $52,763
d. $53,763
e. none of the above

  1. Suppose your annual income is $50,000 and your lender will allow you to have a mortgage payment that is no more than 33% of your monthly income after making other debt payments and paying property taxes, which in your case amount to $500 per month. If the current 30-year mortgage rate is 5%, approximately how large a mortgage can you qualify for (assuming you have the required cash for the necessary down payment)?

a. $160,000
b. $161,000
c. $162,000
d. $163,000
e. none of the above

3. Which of the following borrowers would be considered subprime. One with a FICO score of

a. 650

b. 675
c. 700
d. 750
e. none of above

4. Which of the following purchase mortgages?

I. FNMA   II. FHLMC        III. GNMA


a. I and II only
b. I and III only
c. II and III only
d. all of the above
e. none of the above
Answer #1

(1) Mortgage = $ 1000000, Interest Rate = 5 %, Maturity = 30 years and Balloon Payment = $ 750000 (will come in at the end of Year 30)

Let the annual mortgage payments be $ P

Therefore, 1000000 = P x (1/0.05) x [1-{1/(1.05)^(30)}] + 750000 / (1.05)^(30)

1000000 = 173533.09 + P x (1/0.05) x [1-{1/(1.05)^(30)}]

826466.91 = P x 15.372451

P = 826466.91 / 15.372451 = $ 53762.86 ~ $ 53763
Hence, the correct option is (d)

NOTE: Please raise separate queries for solutions to the remaining unrelated questions, as one query is restricted to the solution of only one complete question with a maximum of four sub-parts.